A Bibliography of Publications of Andrzej Hanyga

Andrzej Hanyga University of Bergen

Institute for Solid Earth Physics

Alle’gaten 41, N5007 Bergen Norway

Tel: ?n/a? FAX: ?n/a?

E-mail: Andrzej.Hanyga@ifjf.uib.no (Internet)

13 October 2017 Version 1.06

Abstract

This bibliography records publications of Andrzej Hanyga.

Title word cross-reference

2D [LLH95].

3D [GH95].

anisotropic [Han84]. Application [HP95a]. Applications [Han93]. Applied [Han94a]. arrivals [Han94c]. Asymptotic [Han89a, Han93, Han94a, Han95, LLH95, GH95].

behaviour [Han89b]. Boundary [Han89a].

case [Han91]. caustic [HS91]. caustics [Han84, Han89b, HS91]. complex [HP95b]. computation [Han84]. curve [HP95b]. curved [HP95a]. cusps [HS91].

Diffraction [Han89a, HS91, Han93, Han94a, Han94b, Han94d, Han95, HP95a]. domain [Han94d, HP95a].

Earth [HLP84]. Edge [Han94a, Han94d, Han95, HP95a]. edge-and-vertex [Han94d, Han95, HP95a]. Edge-Vertex [Han94a]. effects [Han89a]. elastic [Han84].

field [Han94b]. functions [LLH95, GH95].

Generalized [HH90, HH94]. geological [HP95b]. Green [LLH95, GH95].

III [Han89a].

Local [Han89b].

media [Han84]. modelling [HP95a]. models [HP95b]. multiple [Han91, Han94c]. multiple-valued [Han91]. non [HP95a]. non-smooth [HP95a]. Numerical [Han84, Han88, Han93].

Point [HP95b]. Point-to-curve [HP95b]. presence [Han84]. Propagation [HLP84]. Pt [Han89a]. pulses [HS91].

Ray [HH90, Han91, Han94c, HH94, Han94b, HP95b]. rays [Han88]. reflection [HP95a].

Seismic [HLP84, HP95a]. seismograms [HH90, HH94]. simple [Han89b, HS91]. singularities [Han94b]. smooth [HP95a]. surfaces [HP95a]. Synthetic [HH90, HH94].

Theory [Han89a, Han93, Han94a, Han94d, Han95]. Time [Han94d, HP95a]. Time-domain [Han94d, HP95a]. times [Han91]. Tracing [HH90, HH94, Han88, Han91, Han94c, HP95b]. travel [Han91].

valued [Han91]. Vertex [Han94a, Han94d, Han95, HP95a]. vicinity [HS91].

Wave [HLP84]. wavefields [Han84, Han89b]. wavefronts [Han88].

References

green3

[GH95] Paulo Lucio G. Lambaré and A. Hanyga. 3D asymptotic green’s functions. submitted for the Annual EAEG Meeting in Glasgow, 1995.

Hanyga:1984:NCE

[Han84] A. Hanyga. Numerical computation of elastic wavefields in anisotropic elastic media in the presence of caustics. In L. B. Felsen, editor, Hybrid Formulation of Wave Propagation and Scattering. Martinus Nijhoff Pub., Dordrecht, The Netherlands, 1984.

Hanyga:1988:NTR

[Han88] A. Hanyga. Numerical tracing of rays and wavefronts. In D. J. Doornbos, editor, Seismological Algorithms, chapter II.2, pages 169– 233. Academic Press, New York, NY, USA, 1988.

Hanyga:1989:BEA

[Han89a] A. Hanyga. Boundary effects in asymptotic diffraction theory, pt I– III. Seismo-series, 35-37, 1989.

Hanyga:1989:LBW

[Han89b] A. Hanyga. Local behaviour of wavefields at simple caustics. Seismo-series, 38, 1989.

Hanyga:1991:RTC

[Han91] A. Hanyga. Ray tracing in the case of multiple-valued travel times. In Expanded Abstracts. Society of Exploration Geophysicists, Tulsa, OK, 1991.

Hanyga:1993:NAA

[Han93] A. Hanyga. Numerical applications of asymptotic diffraction theory. In R. Kleinman et al., editors, Mathematical and Numerical Aspects of Wave Propagation, pages 246–258. SIAM, Philadelphia, PA, USA, 1993.

Hanyga:1994:ADT

[Han94a] A. Hanyga. Asymptotic diffraction theory applied to edge-vertex diffraction. In S. Rionero and T. Ruggeri, editors, Proceedings of the 7th Conference on Waves and Stability in Continuous Media. World Scientific Publishing Co., Singapore, Philadelphia, River Edge, NJ, 1994.

Hanyga:1994:DAR

[Han94b] A. Hanyga. Diffraction at ray field singularities. In V. Buslaev, editor, Asymptotic methods of the spectral theory. Gordon & Breach, 1994.

Hanyga:1994:RTM

[Han94c] A. Hanyga. Ray tracing of multiple arrivals. submitted to Geophys. J. Int., 1994.

Hanyga:1994:TDT

[Han94d] A. Hanyga. Time-domain theory of edge-and-vertex diffraction. submitted to Nonlinear Processes in Geophysics, 1994.

Hanyga:1995:AEV

[Han95] A. Hanyga. Asymptotic edge-and-vertex diffraction theory. submitted to Geophys. J. Int., 1995.

Hanyga:1990:SSG

[HH90] A. Hanyga and H. B. Helle. Synthetic seismograms from generalized ray tracing. In Expanded Abstracts. Soc. Expl. Geophys., Tulsa, OK, 1990.

Hanyga:1994:SSG

[HH94] A. Hanyga and H. B. Helle. Synthetic seismograms from generalized ray tracing. Geophysical Prospecting, 43:51–76, 1994.

Hanyga:1984:SWP

[HLP84] A. Hanyga, E. Lenartowicz, and J. Pajchel. Seismic Wave Propagation in the Earth. Elsevier Science Publishers, Amsterdam, The Netherlands, 1984.

Hanyga:1995:ATD

[HP95a] A. Hanyga and J. Pajchel. Application of time-domain edge-and-vertex diffraction in seismic modelling of reflection from curved and non-smooth surfaces. 1995.

Hanyga:1995:PCR

[HP95b] Andrzej Hanyga and Jan Pajchel. Point-to-curve ray tracing in complex geological models. Geophysical Prospecting, 43(7):859–872, October 1995. CODEN GPPRAR. ISSN 0016-8025. URL http://

                                        onlinelibrary.wiley.com/journal/
                                        10.1111/%28ISSN%291365-2478.

Hanyga:1991:DPV

[HS91] A. Hanyga and M. Seredyńska. Diffraction of pulses in the vicinity of simple caustics and caustic cusps. Wave Motion, 14:101–121, 1991.

Lambare:1995:AGF

[LLH95] G. Lambaré, Paulo Lucio, and A. Hanyga. 2D asymptotic green’s functions. submitted to Geophys. J. Int., 1995.